Outage : Problem with internet services – DDoS attacks on 28/08-31/08/20 : SOLVED

Dear customers,

We have experienced an outage with internet services (traffic drops) due to several DDoS attacks*** on our network. The attacks occurred at different hours and were causing intermittent connection, speed issues and packet loss. The heaviest one took place on Monday 31/08 and lasted from 11h15 until 18h45. Outside of those hours, everything was normal. Our Network Operations team has found and applied (on 01/09 at 1h29) a permanent solution to prevent and immediately mitigate such attacks in the future.

Please find all details below:

Start*: 2020-08-28 18:00*
End*: 2020-09-01 01:29*
Duration: 1) 2020-08-28 18:00 – 2020-08-28 19:00

2) 2020-08-30 20:30 – 2020-08-30 20:45

3) 2020-08-30 22:20 – 2020-08-31 00:55

4) 2020-08-31 04:05 – 2020-08-31 05:05

5) 2020-08-31 11:15 – 2020-08-31 18:45

Impacted services: all internet services
Affected Areas**: all areas

EDIT 12:30*: We have been under attack again since approximately 11:15. Access to our issues blog was limited so we were unable to keep you updated. Our apologies. We are doing our best to mitigate and solve the problem.

EDIT 2: 15h12*: Seems that the last attack is still ongoing.  We are working hard to try to minimize the impact. Rebooting your modem/router to receive new DNS servers can help (but not guaranteed).  In general surfing should be possible, only certain destinations will be difficult to reach.  We keep you posted.

EDIT 3: 20h25*: The last attack finished at 18:45. We did our best to mitigate what was possible but the overload was extensive. Edpnet DNS servers are currently available again. We are working very hard with an external party towards providing a permanent solution to minimize the impact of any further attacks.

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site

*** DDoS (Distributed-Denial-of-Service- attack) is kind of a cyber-attack, when malicous users seek to make a machine or network resource unavailable by flooding the provider or target with superfluous requests, which results in systems overload and connectivity issues


Outage : Problem with internet services – SOLVED

Dear customers,

We are currently experiencing an outage with internet services (traffic drops). Our Network Operations team is investigating the outage.

Further details on the affected areas and an estimated duration will be released as soon as possible.

EDIT 19:40*: Our DNS-servers were the target of a concerted DDOS-attack for approximately one hour (18:00 – 19:00), involving large amounts of traffic and major network saturation. At this time the attack has cleared up and we are investigating further.

EDIT 30/8 20:35*: It seems like we’re being attacked again since a few minutes. Our Network Operations team is investigating. More updates will follow soon.

EDIT 20:45*: The attack was mitigated after some actions from our Network Operations team. It was smaller than last time, but still very much able to knock out our DNS-servers. We’re working to restore all services.

EDIT: 21:15*: All services are restored.

EDIT: 31/8 5h05 We experienced other waves of DDoS attacks between 30/8 22h20 and 31/08 05h05. For the moment, the attacks are fully mitigated.

We apologize for the inconveniences.

Best regards
The edpnet team


Outage notification: Internet services down – regions: Ath & Châtelet : SOLVED

UPDATE 11h45: The issue is also solved for Ath

UPDATE 11h25: Outage in Châtelet has been solved.

Dear customers,

We are currently experiencing an outage with internet services in regions Ath and Châtelet. Our Network Operations team is investigating the outage.

Further details on the affected areas and an estimated duration will be released as soon as possible.

We apologize for the inconveniences.

Best regards
The edpnet team


Outage: Internet services down : SOLVED

UPDATE 14h15: all links have been restored. The problem is completely resolved. Further details about this major incident will be shared later this week.

UPDATE 13h05: main path is restored , the peerings are being re-enabled , the traffic flow should get back to normal very soon.

Dear customers,

We are currently experiencing a fiber cut within our network. Internet and telephony services are unavailable for our customers in the Netherlands. Our Network Operations team is investigating the outage.

Further details  and an estimated duration will be released as soon as possible.

We apologize for the inconveniences.

Best regards
The edpnet team


Outage: Internet services down : SOLVED

UPDATE 14h15: all links have been restored. The problem is completely resolved. Further details about this major incident will be shared later this week.

UPDATE 13h05: main path is restored , the peerings are being re-enabled , the traffic flow should get back to normal very soon.

UPDATE 11h49: Sorry for the delay in our update.  We are hoping to solve the root cause of our problem as soon as possible.  If you want to read some background information, you can read this Datanews article:

Dutch: https://datanews.knack.be/ict/nieuws/glasvezelprobleem-zorgt-voor-panne-bij-edpnet/article-news-1598649.html
French: https://datanews.levif.be/ict/actualite/un-probleme-de-fibre-optique-provoque-une-panne-chez-edpnet/article-news-1288179.html

UPDATE 10h13: Internet services in all areas are affected.  Our engineers are working on it.

Dear customers,

We are currently experiencing a fiber outage within our network. You may notice capacity issues. Our Network Operations team is investigating the outage.

Further details and an estimated duration will be released as soon as possible.

We apologize for the inconveniences.

Best regards
The edpnet team


Outage: DSL services (regions: Namur/Charleroi/Liège/Mons) #SOLVED

Dear customers,

We are currently experiencing an outage with our DSL services in the southern provinces.  Our Network Operations team is investigating this outage.  Further updates will be released as soon as possible.

Please find all details below:

Start*: 2020-03-31 06:29*
End*: 2020-03-31 11:15*
Duration: 4:45h
Impacted services: All DSL services
Affected Areas**: 41GRE: Liège, Namur, province of Luxembourg
71GIL: Charleroi, Mons, Tournai, Namur

EDIT 10:30*: The cause is found. During a migration to proactively improve our capacity due to the increasing bandwidth usage, an issue with our supplier’s hardware occurred causing them to block traffic. We are currently working with the supplier to rectify this problem as soon as possible, this happens region per region.

EDIT 11:15*: All regions have been brought back online. The issues should be resolved. It is possible you will need to restart your modem to come back online.

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: Telephony capacity issues : SOLVED

Dear customers,

Due to increased usage of communications facilities, you may notice some calls failing. We are now working in cooperation with other operators to get it resolved as soon as possible.  Further updates will be released as soon as possible.

Please find all details below:

Start*: 2020-03-16 09:00
End*: n/a
Duration: n/a
Impacted services: telephony
Affected Areas**: All areas

UPDATE 2020-03-18 10:00h : All telephony operators have done necessary changes. The voice traffic is stable now.

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: Connectivity issues – region : Brussels : SOLVED

Dear customers,

We have experienced an outage with internet connectivity in Brussels region, due to an issue on a fiber. The outage started at 01:15 but was mitigated by the backup path until around 4:00, when the link restored partially. This caused traffic to some areas to be routed over the primary link again, resulting in connection issue for these areas, until it was fully resolved at 7:35.

Please find all details below:

Start*: 2020-01-21 01:15 / 04h04
End*: 2020-01-21 07:35
Duration: 3 hours 21 minutes
Impacted services: All internet services
Affected Areas**: 02STR: Brussels

 

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: Outgoing mails to @icloud.com, @mac.com, @me.com [SOLVED]

Dear customers,

We are currently experiencing issues with mail delivery to certain domains, such as @icloud.com, @mac.com, @me.com, because Apple is currently blocking our outbound mail servers. Please contact Apple support if you have issues with mail delivery from edpnet to these domains. Contacting them will help to unblock as soon as possible.

Start*: 2020-01-17
End*: 2020-02-04
Duration: Approx. 2 1/2 weeks
Impacted services: Mail delivery to @icloud.com, @mac.com, @me.com
Affected Areas**: All areas

EDIT 2020-02-04 9:00h*:  The issue has been resolved and all mail from edpnet mailservers (with correctly configured domains) will be promptly delivered to all Apple mail addresses.

 

Best regards
The edpnet team

 


Outage: problem with internet services : SOLVED

Dear customers,

Due to an issue with one of our upstream providers, a part of the internet was not reachable anymore. We were rerouting the traffic to another provider while fixing the issue. Since Saturday 14/12/2019 8h45 CET the previous configuration is restored and traffic runs its normal path.

Please find all details below:

Start*: 2019-12-13 15:15
End*: 2019-12-13 16:07
Duration: 52 minutes
Impacted services: all internet services
Affected Areas**: Belgium & The Netherlands

 

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site